GRAPH f(x)=-(x+3)+2 whats the Y intercept?
Answer by KMST(5396)

That is a linear function, and the graph will be a straight line. You only need two points to determine a straight line.
You could calculate the value of the function for two different values of x, plot the points, and draw the line.
For example, for x=-3, the value of the function is
f(-3)=-(-3+3)+2=-0+2=2
and you would plot the point A(-3,2) on your graph.
Then if you pick x=4, for x=4, the value of the function is
f(4)= -(4+3)+2=-7+2=-5
and you would plot the point B(4,-5) on your graph.
The y-intercept is the point where the line intersects the y-axis, which is the vertical line described by the equation x=0. So the point with x=0 and y=f(0) is the y-intercept, but usually you just give they value, because everybody knows that x=0 is the x value.
In your case, f(0)=-(0+3)+2=-3+2=-1.